OpenStudy (anonymous):

Area of a circle is 50 inches fnd the radius.

OpenStudy (anonymous):

*93 not 50

OpenStudy (anonymous):

Well the area A = Pi * r^2 Then A / Pi = r^2 \[r = \sqrt{\frac{ A }{ \pi }}\]

OpenStudy (whpalmer4):

Area can't be "50 inches" because the inch is a unit of length, whereas area has dimensions length*length In any case, the area of a circle is given by \[A=\pi r^2\] You know A, you know \(\pi \approx 3.1415926535\), you just solve for \(r\)